OpenStudy (anonymous):

does anyone know how to solve using equivalent equation/expression? im doing a unit solving systems (algebra) my question is -2x+y=8 and theres another equation to this same problem which is 3x+2y=9 i also have to solve using graphs and by substitutions

OpenStudy (anonymous):

solve either of the equations for x or y

OpenStudy (anonymous):

thanks but also when i was done i got y=8=2x

OpenStudy (anonymous):

just solve one for either x or y

OpenStudy (anonymous):

oh ,lol im slow haha thanks

OpenStudy (anonymous):

No my first sentence was too ambiguous

OpenStudy (anonymous):

Need any more help?

OpenStudy (anonymous):

oh its ok and how would i solve it for graphs?

OpenStudy (anonymous):

Well I am not sure exactly what they want

OpenStudy (anonymous):

they want me to graph what i got for the equation

OpenStudy (anonymous):

What did you get for the equation? I think maybe they want you to add the equations and graph the answer

OpenStudy (anonymous):

i got 2x+8=y and y=4.5-1.5x so would i set them equal to eac other to get the eqation for the graph?

OpenStudy (anonymous):

I am not sure. If you set the two equal to each other that would be finding the point at which the two functions intersect

OpenStudy (anonymous):

hmm yeah you right

OpenStudy (anonymous):

Do you have the original question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If you are trying to solve this system of equations then you only need to know the point at which the two graphs intersect because that point gives one x and one y such that both equations are true

OpenStudy (anonymous):

-2x+y=8} 3x+2y=9 } solve by graphing question a solve by using equivelent equations/expressions question 2 yes but how do i get it?

OpenStudy (anonymous):

So, we need to set the two functions equal and to make it easier on our selves lets solve both for y

OpenStudy (anonymous):

ok(:

OpenStudy (anonymous):

so for the first i got y=8+2x

OpenStudy (anonymous):

y = (9-3x)/2 and y = 8 + 2x Nice job So for what x is 8 + 2x = (9-3x)/2?

OpenStudy (anonymous):

and the second y=4.5-1.5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

totally, I wrote mine a little differently but they are equvalent

OpenStudy (anonymous):

Well, we just need to solve for x!

OpenStudy (anonymous):

thankyou so much!

OpenStudy (anonymous):

Did you get a value for x yet?

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

i got- 7

OpenStudy (anonymous):

8 + 2x = (9-3x)/2 Multiply both sides by 2: 2(8 + 2x) = 9-3x, 16 + 4x = 9 - 3x Add 3x to both sides: 16 + 7x = 9 Subtract 16 from both sides: 7x = -7 Divide both sides by 7: x = -1 Maybe check your multiplication and addition on that one

OpenStudy (anonymous):

Now we only need our y coordinate

OpenStudy (anonymous):

so 8+2(-1)=6 so(-1,6)?

OpenStudy (anonymous):

So I am going to pick one of the equations and substitute x = -1 in y = 8 + 2(-1) y = 8 -2 y = 6 So now we have our solution kind of by graphing. We found the point at which the two equations intersect and you are correct it is (-1,6)

OpenStudy (anonymous):

Good luck, keep up the good work

OpenStudy (anonymous):

ok so i put the point on the graph at (-1,6)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Oh I'm sorry is there something else you needed with?

OpenStudy (anonymous):

oh it just that what do i do to plot the other points?

OpenStudy (anonymous):

The two equations or functions whatever you want to call them y= 4.5 - 1.5x and y = 8 + 2x The y-intercept of the first I am going to call the first function f(x) so as not to get confused is 4.5 and the slope of f(x) is -1.5 The y-intercept of the second which ill call g(x) is 8 and its slope in 2

OpenStudy (anonymous):

so i would so two down and two pos. (right) in a linear function

OpenStudy (anonymous):

two up and two right

OpenStudy (anonymous):

ok thank you so much for you time i really appreciate it thank you(:

OpenStudy (anonymous):

No problem it was fun

OpenStudy (anonymous):

(: